OpenStudy (anonymous):

Which of the two functions below has the largest maximum y-value?f(x) = -3x^4 - 14g(x) = -x^3 + 2 A.There is not enough information to determine B.g(x) C.f(x) D.The extreme maximum y-value for both f(x)and g(x) is infinity

OpenStudy (anonymous):

you can differentiate both equation and find the value of x. First equation Dy/dx=-12x^3 The minimum value can be found by dy/dx=0 So -12x^3=0 X=0 Substitute this value on the equation and the result is the value wanted F(0)=0-14 F(0)=-14 Second equation Dy/dx=-3x^2 Equal 0 -3x^2=0 X=0 Substitute on the equation G(0)=0+2 G(0)=2 G(x)>f(x)